Circular No. 6435

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRS 1915+105 B. Thomas, Hughes STX and NASA Goddard Space Flight Center

(GSFC); R. Corbet, Universities Space Research Association and GSFC; J. Swank, GSFC; W. Focke, University of Maryland and GSFC; and F. Marshall, GSFC, report: “Observations made of GRS 1915+105 on July 11.357–11.397 UT with the RXTE PCA show the presence of strong x-ray (1–12 keV) quasi-periodic oscillations (QPOs) at 2.64 Hz (rms amplitude 9.2 ± 0.3 percent, Lorentzian width 0.25 Hz) with a weaker feature at 5.27 Hz (3.3 ± 0.4 percent, width 2 Hz). We note that the previously observed QPO between 0.5 and 100 mHz and at 67 Hz (IAUC 6392) were not apparent in this observation. This change in properties may be related to the recent radio outburst from this source (IAUC 6432). The flux during this observation was ∼ 700 mCrab and the variability < 10 percent during the observation interval.”

COMET C/1996 N1 (BREWINGTON) Improved ephemeris from orbital elements on MPEC 1996-N02:

1996 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
July 6 11^{h}38.^{m}78 + 5^{o}49.^{′}3 0.833 1.053 68.^{o}6 64.^{o}0  9.8
11 11 53.46 +10 38.6 0.837 1.015 65.5 65.8  9.7
16 12 08.40 +15 21.1 0.845 0.982 62.9 67.1  9.6
21 12 23.50 +19 54.8 0.855 0.956 60.7 68.0  9.5
26 12 38.72 +24 17.8 0.867 0.938 59.1 68.3  9.4
31 12 54.03 +28 28.4 0.881 0.928 58.1 68.2  9.4
Aug. 5 13 09.49 +32 25.6 0.895 0.926 57.6 67.7  9.4
10 13 25.21 +36 08.4 0.908 0.933 57.8 66.8  9.5
15 13 41.39 +39 36.4 0.921 0.948 58.5 65.6  9.6
20 13 58.30 +42 49.5 0.933 0.972 59.8 64.1  9.7
25 14 16.28 +45 47.8 0.944 1.002 61.6 62.5  9.9
30 14 35.71 +48 31.1 0.953 1.038 63.8 60.8 10.1
Sept. 4 14 57.04 +50 58.8 0.960 1.080 66.5 58.9 10.2

Visual m_{1} and coma-diameter estimates: July 10.91 UT, 9.7, – (A. Pereira, Cabo da Roca, Portugal, 0.25-m reflector); 12.09, 9.4, 2′.5 (J. E. Bortle, Stormville, NY, 0.41-m reflector); 12.85, 8.8, 8′ (A. Baransky, Kiev, Ukraine, 20×110 binoculars); 14.22, 9.0, 8′.5 (C. S. Morris, Lockwood Valley, CA, 20×80 binoculars).

1996 July 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green